Bodacious Bags buys nails in 15-pound boxes but their growth has increased their need to an average of 3870 boxes a year. Preparing an order and receiving a shipment of nails involves a cost of $1.90 per order. Annual carrying costs are $0.5 per box.

What will be their total cost of ordering and carrying the nails? (Include the pennies in your answer)

Your Answer:
Answer #1

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST

ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST

EOQ = SQRT(2 * 3780 * 1.9 / 0.5) = 169.49

ANNUAL HOLDING COST = (169.49 / 2) * 0.5 = 42.37

ANNUAL ORDERING COST = (3780 / 169.49) * 1.9 = 42.37

TOTAL COST = AHC + AOC = 42.37 + 42.37 = 84.74